OpenStudy (cookiemonster2412):

Can someone please help. Which of the following equations has both 1 and -3 as solutions? A. x^2-2x-3=0 B. x^2+2x-3=0 C. x^2-4x+3=0 D. x^2+4x+3=0

OpenStudy (cookiemonster2412):

@acxbox22

OpenStudy (acxbox22):

factor each equation set equal to zero solve for x for the first one... x^2-2x-3=0 (x-3)(x+1)=0 x-3=0 or x+1=0 x=3 or x=-1